Grade 12Physical Chemistry

The density of water of a lake is 1.25 g M/L and one kg of this Easte contains 92 g of Na ions. What is the molarity and molality of Na ions in the water of the lake?

Santoshi Iyer
7 Years agoGrade 12

1 Answer

Shubham Tandia
7 Years ago
Number of moles of Na+ ions = 92/23 = 4 mole
Now Molality= moles of Na/ 1 kg 
= 4/1 = 4m ,Ans.
Now Volm. Of solution=mass of solution/density of sol
=1000g (1kg given) / 1.25 gmL-1
= 800 ml = 0.8 L
So molarity M = moles/litre = 4 mol / 0.8 L = 5M , Ans